OpenStudy (anonymous):

WILL give medal and follow, Meg has a can that contains 80% of orange juice and the rest water. The can has 1 liter of water. Part A: write an equation in one variable that can. Be used to find the toatal number of liters of orange juice and water in the can. Define the variable used in the equation. Part B: how many liters of orange juice are present in the can? SHOW YOUR WORK.

OpenStudy (mosaic):

Let "x" be the total number of liters of orange juice and water combined. 80% of the mixture is pure orange juice which means 20% of it is water. Therefore, x * 20/100 = 1 is the equation in single variable that can be used to find x.

OpenStudy (mosaic):

Solve the above equation for x. x * 20/100 = 1 x * 1/5 = 1 x = 5. So there is a total of 5 liters of orange juice and water mixture. 1 L of it is water. Therefore, the rest of 5-1 = 4 liters is orange juice present in the can.

OpenStudy (anonymous):

@mosaic

OpenStudy (anonymous):

U saved me